# Online Intrinsic Rewards for Decision Making Agents from Large Language Model Feedback

### v2

@@ Abstract
Automatically synthesizing dense rewards from natural language descriptions is a promising paradigm in reinforcement learning (RL), with applications to sparse reward problems, open-ended exploration, and hierarchical skill design. Recent works have made promising steps by exploiting the prior knowledge of large language models (LLMs). However, these approaches suffer from important limitations: they are either not scalable to problems requiring billions of environment samples, due to requiring LLM annotations for each observation, or they require a diverse offline dataset, which may not exist or be impossible to collect. In this work, we address these limitations through a combination of algorithmic and systems-level contributions. We propose ONI, a distributed architecture that simultaneously learns an RL policy and an intrinsic reward function using LLM feedback. Our approach annotates the agent’s collected experience via an asynchronous LLM server, which is then distilled into an intrinsic reward model. We explore a range of algorithmic choices for reward modeling with varying complexity, including hashing, classification, and ranking models. By studying their relative tradeoffs, we shed light on questions regarding intrinsic reward design for sparse reward problems. Our approach achieves state-of-the-art performance across a range of challenging, sparse reward tasks from the NetHack Learning Environment in a simple unified process, solely using the agent’s gathered experience, without requiring external datasets. We make our code available at https://github.com/facebookresearch/oni.
@@ 1 Introduction
  1. 1. Generating the reward function’s code by LLM. A number of methods have been proposed to automatically generate executable code that computes the reward directly (Ma et al., 2023; Xie et al., 2023; Yu et al., 2023; Li et al., 2024). While they have demonstrated success in complex continuous control tasks, they either require access to environment source code to include in the prompt, or a detailed description of input parameters and reward function templates. Furthermore, they are limited to reward functions compactly expressible via code, describing explicit logic; and it is unclear how these approaches can easily process high-dimensional state representations such as images, or semantic features such as natural language.
  2. 2. Generating reward values by LLMs. Motif (Klissarov et al., 2023) is a typical example of this category. It ranks the captions of pairs of observations using an LLM and distills these preferences into a parametric reward model. Motif does not require access to environment source code nor numerical state representation, can process semantic input features, and can scale to problems requiring billions of environment samples. Nevertheless, it also suffers from two important limitations. First, it requires a diverse, pre-existing dataset of captioned observations which are used to elicit preferences from the LLM. In many situations, such a dataset might not exist, and collecting it can increase the sample complexity. More importantly, collecting a diverse dataset often requires a non-trivial reward function that is feasible to optimize, which is the primary problem we aim to solve with intrinsic reward functions in the first place. Second, it involves a complex three-stage process, which sequentially annotates observations using an LLM, trains a reward model, and finally trains an RL agent. This is still time-consuming, given that the LLM annotation process can take several days’ worth of GPU hours, and is done prior to training the reward model and RL agent. Alternatively, Chu et al. (2023) query the LLM to directly label observations as having high or low reward at each timestep. However, querying an LLM for every observation is computationally infeasible for many RL applications, which involve millions or billions of observations.
@@
In this work, we present ONI, a distributed online intrinsic reward and agent learning system. ONI assumes access to captions of observations, similar to previous work Klissarov et al. (2023); Chu et al. (2023). The captions of collected observations are annotated by an asynchronous LLM server, and both the policy and intrinsic reward model are simultaneously updated using LLM feedback. ONI removes the dependency on external datasets beyond the agent’s own experience and enables large-scale RL training with ease. Such a learning framework allows us to systematically compare different algorithmic choices for synthesizing LLM feedback. Specifically, we explore three methods: the first one is retrieval-based and simply hashes the annotations; the second builds a binary classification model to distill the sentiment labels returned by the LLM; and the third sends pairs of captions to the LLM server for preference labeling and learns a ranking model, similar to Motif. By carefully comparing the three proposed algorithms, we provide valuable insights into several important questions regarding intrinsic reward design. We demonstrate that ONI is able to match Motif’s performance across a range of challenging, sparse rewards from the NetHack Learning Environment (NLE) (Küttler et al., 2020), solely using the agent’s gathered experience in a single, unified process.
@@ 2 Background
We consider a partially observed Markov decision process (POMDP) setting where the problem is defined by $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{O},p_{0},P,O,r,\gamma)$. At each episode, an initial state $s_{0}\in\mathcal{S}$ is sampled from the initial state distribution $p_{0}$. At each time step $t$, the agents observes $o_{t}\in\mathcal{O}$ which is computed by the emission function $O(s_{t})$, and takes an action $a_{t}\in\mathcal{A}$. This action causes the environment to transition to a new state, $s_{t+1}\sim p(s_{t},a_{t})$. A new observation $o_{t+1}$ and a reward $r_{t+1}=r(o_{t+1})$ is given to the agent, and the process continues. The goal of the agent is to learn a policy $\pi:\mathcal{O}\rightarrow\Delta(\mathcal{A})$ which maximizes the expected return $\mathbb{E}_{\pi}\left[\sum_{t}\gamma^{t}r_{t}\right]$. In this work, we additionally assume each observation $o_{t}$ includes a textual caption $c_{t}$, which could be empty. For observation spaces without textual captions, $c_{t}$ could in principle be provided by a captioning model as well.
@@
A key research question is how to define or learn the intrinsic reward $r^{\mathrm{int}}$. A first option is to manually define $r^{\mathrm{int}}$ based on task-specific goals, for example a measure of the distance between the agent’s current state and the goal state. However, handcrafting the intrinsic reward function can require significant domain knowledge and must be redone for each new task. A second option is to define $r^{\mathrm{int}}$ to measure some notion of observation novelty, which encourages the agent to systematically explore the environment. This can work well in smaller environments, but fails in ones that cannot be exhaustively explored in a tractable amount of time. A third class of methods, which we focus on in this work, leverage LLMs to automatically synthesize $r^{\mathrm{int}}$ to reflect prior knowledge about the task. We discuss all three classes of methods in Section 4.
@@ 3.1 System design: distributed PPO with LLM annotations
This section outlines the system we have built to learn online intrinsic rewards alongside the policy optimization. The engineering and design here is an important piece of our research, as the rest of our experimental studies are conducted within this system and influenced by the throughput of its interacting components.
@@
Our core system illustrated in Figure 3.1 is built on top of the Sample Factory library v1.0 (Petrenko et al., 2020) and their asynchronous variant of proximal policy optimization (Schulman et al., 2017), referred to as APPO. APPO operates on a single machine and concurrently runs many environment instances while asynchronously updating policy and value estimates with the usual PPO rules, and adequately handles policy staleness and data transfers between these components. Concretely for NetHack, by running 480 environment instances APPO collects approximately 32k environment interactions per second on a Tesla A100-80GB GPU with 48 CPUs.
@@
Figure 3.1: Overall system diagram of ONI. Our additions to Sample Factory are highlighted in red. We added an asynchronously executing LLM server and learned reward function, and connect them back into the main learning process in a way that does not hurt the overall throughput of the policy and value learning.
@@
For online reward learning via LLM annotations in this system111 The prior work by Klissarov et al. (2023) in the offline setting was able to connect their intrinsic reward function into Sample Factory’s APPO implementation with a few lines of code, as it just needed to load the PyTorch module of the learned reward function., we added 1) an LLM server hosted on a separate node, 2) an asynchronously running process that passes observation captions to the LLM server via HTTP request, 3) a hash table that stores the captions and corresponding LLM annotations, and 4) and learning code that dynamically updates a reward model that learns on these annotations. Without being asynchronous, these components have the potential to block the concurrent execution and can reduce the throughput of the overall system, because calling into an LLM and updating the reward model are time-consuming. We added them in a way that retains most of the throughput (!), approximately 80-95% of the original: the average throughput is 30k environment interactions per second if we do not train any additional reward model to distill the LLM annotations, and 26k if a classification-based reward model is learned at the same time (see Section 3.2 for the description of those approaches). The throughput of LLM annotation depends on the actual LLM and prompt we are using. For the reader’s reference, when hosting a LLaMA-3.1-8B server on a Tesla A100-80GB GPU using the prompt displayed in Appendix B, ONI annotates approximately 810k NetHack captions over the whole training process (2B environment steps, 19-20 hours). The rest of this section overviews the relevant components and how we have modified them.
@@
Background: APPO’s Distributed Execution and Shared Memory
APPO has three main types of workers:
@@
Our New Asynchronous LLM Worker and Remote LLM Server
This worker has input and output queues that communicate back with the learner process, which we use so it does not block the main execution of the system. The LLM worker awaits new observations to label from the input queue, formats them into prompts, calls into an LLM, and returns the annotations in the output queue. Additionally, the prompt templates and other LLM options are configured by the LLM worker. We use an annotation and message format that support all the intrinsic reward labels that we consider in Section 3.2. As Sample Factory already utilizes most of the free CPU and GPU capacity of the system, we opted to call into the LLM via an HTTP/REST interface rather than loading it in this process directly. The communication here adds minimal overhead, and also makes us not need to coordinate the shared memory of multiple GPUs between the main APPO code and LLM. Alongside every APPO run, we allocate an unloaded machine with a new instance of VLLM (Kwon et al., 2023b) to exclusively serve the queries from the RL run. Our system also allows using the same node for both processes if computational resources allow, which further reduces communication costs.
One important design decision is that the asynchronous LLM server is only able to process a small percentage of the overall observations encountered 222often $\leq 0.04\%$ for a LLaMA-3.1-8B instance using a single A100-80GB GPU, or $\leq 0.02\%$ when using a V100-32GB GPU, because the environment rollout workers have a much higher throughput than the LLM. An alternative design would be to block the main APPO components and wait for the LLM to label more messages, but we find it more realistic to continue running the policy and label messages as the LLM throughput allows. This also creates the problem of needing to decide on what messages to send to the LLM: the last-in-first-out queue (LIFO), or some uncertainty-based selection. For this work, we use LIFO for simplicity, but note that it would be interesting to investigate alternate approaches in future work.
Our Modified Learner
Lastly, we connect this new LLM worker back into the main learner, and dynamically learn a reward model on the annotations obtained from it. To do this, we modified the two threads of the learner worker:
@@
To prevent overfitting the limited amount of annotations in the early stage of training, the training thread only continuously updates the reward model after we receive $25$k annotations. Before that, the feedback processing thread run a few updates of the reward model every time we receive new annotations. Now that we have a flexible way of annotating messages and fitting a reward model on top of them, we turn to defining the types of intrinsic reward functions.
@@ 3.2 Intrinsic Reward Functions
Retrieval
The simplest approach we consider uses binary labeling and retrieval. The LLM is asked to assign a binary label $y_{i}\in\{0,1\}$ indicating whether a caption $c_{i}$ is “helpful” or “unhelpful” for making progress on the task. The learner worker maintains a hash table $\mathcal{H}$ to store labeled pairs $(c_{i},y_{i})$, managed by the feedback processing thread. In the training thread, each time the RL agent receives an observation $o_{t}$ with caption $c_{t}$, we check if $c_{t}\in\mathcal{H}$ and the intrinsic reward is defined as:
@@
Classification
The second approach we consider is based on binary labeling together with training a classification model. Similarly to above, we label observation captions $c_{i}$ with labels $y_{i}\in\{0,1\}$ indicating whether they are helpful or unhelpful via LLM. We simultaneously train a binary classification model to predict $y_{i}$ from $o_{i}$. More precisely, we model $P(y=1|o)$ by $r^{\mathrm{int}}_{\phi}:\mathcal{O}\rightarrow[0,1]$, which is then used to compute the binary intrinsic reward by thresholding it at $\eta$:
@@
where $\mathbb{I}$ is the indicator function. We study the impact of $\eta$ in Section 5.3. Unlike the previous approach, this method has potential to generalize to observations whose captions are similar, but not identical, to the captions labeled by the LLM. However, like the previous approach, it will assign a same reward to observations which are slightly positive (such as finding a few gold pieces in NetHack) and very positive (finding hundreds of gold pieces or a rare artifact).
Ranking
The third approach we consider is based on ranking observations via pairwise classification, which is the approach taken by Motif. Here, pairs of observations $(o_{1},o_{2})$ are sent to the LLM, which returns a preference label $y\in\{1,2,\varnothing\}$ indicating whether $o_{1}$ or $o_{2}$ is more desirable for accomplishing the task, or if they are equivalent. A reward model $r^{\mathrm{int}}_{\phi}:\mathcal{O}\rightarrow\mathbb{R}$ is trained by minimizing the negative log-likelihood :
@@
where we use average log-likelihood when $y=\varnothing$ [^3] In the equivalent cross entropy minimization formulation, this amounts to using a uniform target $[\frac{1}{2},\frac{1}{2}]$ when $y=\varnothing$. and use the Bradley-Terry model Bradley & Terry (1952) $P_{\phi}(o_{1}\succ o_{2})=1-P_{\phi}(o_{1}\prec o_{2})=\exp\big{(}r^{\mathrm{int}}_{\phi}(o_{1})\big{)}/\big{[}\exp\big{(}{r^{\mathrm{int}}_{\phi}(o_{1})}\big{)}+\exp\big{(}r^{\mathrm{int}}_{\phi}(o_{2})\big{)}\big{]}$. Motif computes the mean $\mu_{\mathcal{D}}$, standard deviation $\sigma_{\mathcal{D}}$, and a fixed quantile $\nu_{\mathcal{D}}$ of $r^{\mathrm{int}}_{\phi}$ over the offline dataset of annotations $\mathcal{D}$. During RL training, it normalizes and thresholds $r^{\mathrm{int}}_{\phi}$ to give the reward

### v3

@@ Abstract
Automatically synthesizing dense rewards from natural language descriptions is a promising paradigm in reinforcement learning (RL), with applications to sparse reward problems, open-ended exploration, and hierarchical skill design. Recent works have made promising steps by exploiting the prior knowledge of large language models (LLMs). However, these approaches suffer from important limitations: they are either not scalable to problems requiring billions of environment samples, due to requiring LLM annotations for each observation, or they require a diverse offline dataset, which may not exist or be impossible to collect. In this work, we address these limitations through a combination of algorithmic and systems-level contributions. We propose ONI, a distributed architecture that simultaneously learns an RL policy and an intrinsic reward function using LLM feedback. Our approach annotates the agent’s collected experience via an asynchronous LLM server, which is then distilled into an intrinsic reward model. We explore a range of algorithmic choices for reward modeling with varying complexity, including hashing, classification, and ranking models. Our approach achieves state-of-the-art performance across a range of challenging tasks from the NetHack Learning Environment, while removing the need for large offline datasets required by prior work. We make our code available at https://github.com/facebookresearch/oni.
@@ 1 Introduction
  1. 1. Generating the reward function’s code by LLM. A number of methods have been proposed to automatically generate executable code that computes the reward directly (Ma et al., 2023; Xie et al., 2023; Yu et al., 2023; Li et al., 2024). While they have demonstrated success in complex continuous control tasks, they either require access to environment source code to include in the prompt, or a detailed description of input parameters and reward function templates. Furthermore, they are limited to reward functions compactly expressible via code, describing explicit logic, and it is unclear how these approaches can easily process high-dimensional state representations such as images, or semantic features such as natural language.
  2. 2. Generating reward values by LLMs. Motif (Klissarov et al., 2023; 2025) is a typical example of this category. It ranks the captions of pairs of observations using an LLM and distills these preferences into a parametric reward model. Motif does not require access to environment source code nor numerical state representation, can process semantic input features, and can scale to problems requiring billions of environment samples. Nevertheless, it also suffers from two important limitations. First, it requires a diverse, pre-existing dataset of captioned observations which are used to elicit preferences from the LLM. In many situations, such a dataset might not exist, and collecting it can increase the sample complexity. More importantly, collecting a diverse dataset often requires a non-trivial reward function that is feasible to optimize, which is the primary problem we aim to solve with intrinsic reward functions in the first place. Second, it involves a complex three-stage process, which sequentially annotates observations using an LLM, trains a reward model, and finally trains an RL agent. This is still time-consuming, given that the LLM annotation process can take several days’ worth of GPU hours, and is done prior to training the reward model and RL agent. Alternatively, Chu et al. (2023) query the LLM to directly label observations as having high or low reward at each timestep. However, querying an LLM for every observation is computationally infeasible for many RL applications, which involve millions or billions of observations.
@@
In this work, we present ONI, a distributed online intrinsic reward and agent learning system — Figure 1 overviews the high-level components. ONI removes the dependency on external datasets beyond the agent’s own experience and enables large-scale RL training with ease. ONI assumes access to captions of observations, similar to previous work Klissarov et al. (2023); Chu et al. (2023). The captions of collected observations are annotated online by an asynchronous LLM server, and both the policy and intrinsic reward model are simultaneously updated using the LLM’s feedback. Such a learning framework allows us to easily instantiate different algorithmic choices for synthesizing LLM feedback. Specifically, we explore three methods: the first one is retrieval-based and simply hashes the annotations; the second builds a binary classification model to distill the sentiment labels returned by the LLM; and the third sends pairs of captions to the LLM server for preference labeling and learns a ranking model, similar to Motif. We demonstrate that ONI is able to match Motif’s performance across a range of challenging, sparse rewards from the NetHack Learning Environment (NLE) (Küttler et al., 2020), solely using the agent’s gathered experience in a single, unified process. We also open-source our code to facilitate future work on LLM-based intrinsic motivation in settings where large offline datasets are not available.
Figure 1: Overview of Motif (Klissarov et al., 2023) and ONI.
@@ 2 Background
We consider a partially observed Markov decision process (POMDP) setting where the problem is defined by $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{O},p_{0},P,O,r,\gamma)$. At each episode, an initial state $s_{0}\in\mathcal{S}$ is sampled from the initial state distribution $p_{0}$. At each time step $t$, the agent observes $o_{t}\in\mathcal{O}$ which is computed by the emission function $O(s_{t})$, and takes an action $a_{t}\in\mathcal{A}$. This action causes the environment to transition to a new state, $s_{t+1}\sim p(s_{t},a_{t})$. A new observation $o_{t+1}$ and reward $r_{t+1}=r(o_{t+1})$ are given to the agent, and the process continues. The goal of the agent is to learn a policy $\pi:\mathcal{O}\rightarrow\Delta(\mathcal{A})$ which maximizes the expected return $\mathbb{E}_{\pi}\left[\sum_{t}\gamma^{t}r_{t}\right]$. In this work, we additionally assume each observation $o_{t}$ includes a textual caption $c_{t}$, which could be empty. For observation spaces without textual captions, $c_{t}$ could in principle be provided by a captioning model as well.
@@
A key research question is how to define or learn the intrinsic reward $r^{\mathrm{int}}$. A first option is to manually define $r^{\mathrm{int}}$ based on task-specific goals, for example a measure of the distance between the agent’s current state and the goal state. However, handcrafting the intrinsic reward function can require significant domain knowledge and must be redone for each new task. A second option is to define $r^{\mathrm{int}}$ to measure some notion of observation novelty, which encourages the agent to systematically explore the environment. This can work well in smaller environments, but fails in ones that cannot be exhaustively explored in a tractable amount of time. A third class of methods, which we focus on in this work, leverage LLMs to automatically synthesize $r^{\mathrm{int}}$ to reflect prior knowledge about the task. We discuss all three classes of methods in more detail in Section 4.
@@ 3.1 System design: distributed PPO with LLM annotations
This section outlines our system for learning online intrinsic rewards alongside the policy optimization. The engineering and design are an important piece of our research, as they enable scaling to the high-throughput settings we are interested in. The rest of our experimental studies are conducted within this system and influenced by the throughput of its interacting components.
@@
Our core system, illustrated in Figure 2, is built on top of the Sample Factory library v1.0 (Petrenko et al., 2020) and their asynchronous variant of proximal policy optimization (Schulman et al., 2017), referred to as APPO. APPO operates on a single machine and concurrently runs many environment instances while asynchronously updating policy and value estimates with the usual PPO rules, and adequately handles policy staleness and data transfers between these components. Concretely for NetHack, by running 480 environment instances APPO collects approximately 32k environment interactions per second on a Tesla A100-80GB GPU with 48 CPUs.
@@
Figure 2: Overall system diagram of ONI. Our additions to Sample Factory are highlighted in blue. We added an asynchronously executing LLM server and learned reward function, and connect them back into the main learning process in a way that does not hurt the overall throughput of the policy and value learning.
@@
For online reward learning via LLM annotations in this system111 The prior work by Klissarov et al. (2023) in the offline setting was able to connect their intrinsic reward function into Sample Factory’s APPO implementation with a few lines of code, as it just needed to load the PyTorch module of the learned reward function., we added 1) an LLM server hosted on a separate node, 2) an asynchronously running process that passes observation captions to the LLM server via HTTP request, 3) a hash table that stores the captions and corresponding LLM annotations, and 4) learning code that dynamically updates a reward model that learns on these annotations. Without being asynchronous, these components have the potential to block the concurrent execution and can reduce the throughput of the overall system, because calling into an LLM and updating the reward model are time-consuming. We added them in a way that retains most of the throughput (!), approximately 80-95% of the original: the average throughput is 30k environment interactions per second if we do not train any additional reward model to distill the LLM annotations, and 26k if a classification-based reward model is learned at the same time (see Section 3.2 for the description of those approaches). The throughput of LLM annotation depends on the actual LLM and prompt we are using. For the reader’s reference, when hosting a LLaMA-3.1-8B server on a Tesla A100-80GB GPU using the prompt displayed in Appendix B, ONI annotates approximately 810k NetHack captions over the whole training process (2B environment steps, 19-20 hours). The rest of this section overviews the relevant components and how we have modified them.
@@
Background: APPO’s Distributed Execution and Shared Memory APPO has three main types of workers:
@@
Our New Asynchronous LLM Worker and Remote LLM Server This worker has input and output queues that communicate back with the learner process, which we use in a way that does not block the main execution of the system. The LLM worker awaits new observations to label from the input queue, formats them into prompts, calls into an LLM, and returns the annotations in the output queue. Additionally, the prompt templates and other LLM options are configured by the LLM worker. We use an annotation and message format that support all the intrinsic reward labels that we consider in Section 3.2. As Sample Factory already utilizes most of the free CPU and GPU capacity of the system, we opted to call into the LLM via an HTTP/REST interface rather than loading it in this process directly. The communication here adds minimal overhead, and also avoids the need to coordinate the shared memory of multiple GPUs between the main APPO code and LLM. Alongside every APPO run, we allocate an unloaded machine with a new instance of vLLM (Kwon et al., 2023b) to exclusively serve the queries from the RL run. Our system also allows using the same node for both processes if computational resources allow, which further reduces communication costs.
One important design decision is that the asynchronous LLM server is only able to process a small percentage of the overall observations encountered 222often $\leq 0.04\%$ for a LLaMA-3.1-8B instance using a single A100-80GB GPU, or $\leq 0.02\%$ when using a V100-32GB GPU, because the environment rollout workers have a much higher throughput than the LLM. An alternative design would be to block the main APPO components and wait for the LLM to label more messages, but we find it more realistic to continue running the policy and label messages as the LLM throughput allows. This also creates the problem of needing to decide on what messages to send to the LLM: a last-in-first-out queue (LIFO), or some uncertainty-based selection. For this work, we use LIFO for simplicity, but note that it would be interesting to investigate alternate approaches in future work.
Our Modified Learner Lastly, we connect this new LLM worker back into the main learner, and dynamically learn a reward model on the annotations obtained from it. To do this, we modified the two threads of the learner worker:
@@
To prevent overfitting the limited amount of annotations in the early stage of training, the training thread only continuously updates the reward model after we receive $25$k annotations. Before that, the feedback processing thread runs a few updates of the reward model every time we receive new annotations. Now that we have a flexible way of annotating messages and fitting a reward model on top of them, we turn to defining the types of intrinsic reward functions.
@@ 3.2 Intrinsic Reward Functions
Retrieval The simplest approach we consider uses binary labeling and retrieval. The LLM is asked to assign a binary label $y_{i}\in\{0,1\}$ indicating whether a caption $c_{i}$ is “helpful” or “unhelpful” for making progress on the task. The learner worker maintains a hash table $\mathcal{H}$ to store labeled pairs $(c_{i},y_{i})$, managed by the feedback processing thread. In the training thread, each time the RL agent receives an observation $o_{t}$ with caption $c_{t}$, we check if $c_{t}\in\mathcal{H}$ and the intrinsic reward is defined as:
@@
Classification The second approach we consider is based on binary labeling together with training a classification model. Similarly to above, we label observation captions $c_{i}$ with labels $y_{i}\in\{0,1\}$ indicating whether they are helpful or unhelpful via LLM. We simultaneously train a binary classification model to predict $y_{i}$ from $o_{i}$. More precisely, we model $P(y=1|o)$ by $r^{\mathrm{int}}_{\phi}:\mathcal{O}\rightarrow[0,1]$, which is then used to compute the binary intrinsic reward by thresholding it at $\eta$:
@@
where $\mathbb{I}$ is the indicator function. We study the impact of $\eta$ in Section 5.3. Unlike the previous approach, this method has the potential to generalize to observations whose captions are similar, but not identical, to the captions labeled by the LLM. However, like the previous approach, it will assign a same reward to observations which are slightly positive (such as finding a few gold pieces in NetHack) and very positive (finding hundreds of gold pieces or a rare artifact).
Ranking The third approach we consider is based on ranking observations via pairwise classification, which is the approach taken by Motif. Here, pairs of observations $(o_{1},o_{2})$ are sent to the LLM, which returns a preference label $y\in\{1,2,\varnothing\}$ indicating whether $o_{1}$ or $o_{2}$ is more desirable for accomplishing the task, or if they are equivalent. A reward model $r^{\mathrm{int}}_{\phi}:\mathcal{O}\rightarrow\mathbb{R}$ is trained by minimizing the negative log-likelihood :
@@
where we use average log-likelihood when $y=\varnothing$ [^3] In the equivalent cross entropy minimization formulation, this amounts to using a uniform target $[\frac{1}{2},\frac{1}{2}]$ when $y=\varnothing$. and use the Bradley-Terry model Bradley & Terry (1952) $P_{\phi}(o_{1}\succ o_{2})=\exp\big{(}r^{\mathrm{int}}_{\phi}(o_{1})\big{)}/\big{[}\exp\big{(}{r^{\mathrm{int}}_{\phi}(o_{1})}\big{)}+\exp\big{(}r^{\mathrm{int}}_{\phi}(o_{2})\big{)}\big{]}$. Motif computes the mean $\mu_{\mathcal{D}}$, standard deviation $\sigma_{\mathcal{D}}$, and a fixed quantile $\nu_{\mathcal{D}}$ of $r^{\mathrm{int}}_{\phi}$ over the offline dataset of annotations $\mathcal{D}$. During RL training, it normalizes and thresholds $r^{\mathrm{int}}_{\phi}$ to give the reward